OpenStudy (anonymous):

If you put $1,000 in a savings account that pays 5% interest compounded continuously, how much money will you have in your account in 10 years? Assume you make no additional deposits or withdrawals. $1,628.89 $1,051.27 $1,648.72 $5,525.85

OpenStudy (anonymous):

Continuous compound formula is \[P \times e ^{r \times t} \] So If we plug in our values, we should get \[1000 \times e ^{0.05 \times 10}\] Which is equal to appx. 1648.72